OpenStudy (anonymous):

If x=3t and y=5t-2, what is the direct algebraic relationship/equation?

OpenStudy (danjs):

You are looking to find out y, in terms of x? y(x)

OpenStudy (danjs):

solve both for t =.. then you can set them equal, and solve for what you want

OpenStudy (anonymous):

So.. for the first one it would be t=x/3?

OpenStudy (danjs):

yes

OpenStudy (danjs):

second is t = (y+2)/5 set both t's equal since they are the same, then you eliminate the t variable,

OpenStudy (danjs):

\[\frac{ y+2 }{ 5 }=\frac{ x }{ 3 }\] solve for whatever you need

OpenStudy (anonymous):

thanks

OpenStudy (danjs):

welcome